# Online Difficulty Filtering for Reasoning Oriented Reinforcement Learning

### v1

@@ 4.3 Case 2: Learnability in soft prompts
$\begin{split}\mathbb{E}_{\mathbf{y}\sim\pi_{\mathrm{init}}(\cdot|\mathbf{x})}\left[\log\frac{\pi^{*}(\mathbf{y}|\mathbf{x})}{\pi_{\mathrm{init}}(\mathbf{y}|\mathbf{x})}\right]&=\frac{p(\mathbf{x})}{\beta}-\log\left((1-p(\mathbf{x}))+p(\mathbf{x})\exp\left(\frac{1}{\beta}\right)\right)\\ &\leq-\frac{1}{2\beta^{2}}p(\mathbf{x})(1-p(\mathbf{x})).\end{split}$ (14)
Here, RHS is proportional to the variance of $\mathrm{Bernoulli}(p(\mathbf{x}))$. Thus, the reward variance determines the lower bound of the divergence between $\pi_{\mathrm{init}}$ and $\pi^{*}$ given the prompt $\mathbf{x}$:
$\mathbb{D}_{\mathrm{KL}}\left(\pi_{\mathrm{init}}(\mathbf{y}|\mathbf{x})\,\|\,\pi^{*}(\mathbf{y}|\mathbf{x})\right)\geq\frac{p(\mathbf{x})(1-p(\mathbf{x}))}{2\beta^{2}},$ (15)
supporting that the prompts with $p(\mathbf{x})\simeq 0.5$ have the largest learnability.
Hence, soft-hard ($p(\mathbf{x_{\mathrm{hard}}})=\epsilon$) and soft-easy ($p(\mathbf{x_{\mathrm{easy}}})=1-\epsilon$) prompts are expected to provide marginal learnability, and intermediate prompts ($\epsilon\leq p(\mathbf{x})\leq 1-\epsilon$) provides the strongest learning signal. See Appendix B for the full derivation.
@@
1:Initial policy model $\pi_{\mathrm{init}}$; Reward $r$; Prompts queue $\mathcal{Q}$; Pass rate thresholds $T_{\text{Low}},T_{\text{High}}$; Batch size $N$; Group size $G$; $r_{\text{acc}}$ (5); Visit count $\mathrm{vc}(\mathbf{x})$.
2:$\mathcal{P}_{\text{active}}$: The set of examples currently undergoing asynchronous rollout.
3:$C_{\text{max}}$: The maximum number of examples that can be processed concurrently.
@@
21:         Compute $\hat{A}_{i}$ for $\mathbf{y}_{i}$ in $\mathcal{B}^{(t)}$ through group relative advantage estimation (4).
@@ 4.4 Method: online difficulty filtering with fixed batch size
We analyze an online difficulty filtering approach that ensures a fixed batch size throughout training for a reasoning-oriented agent. Unlike static curricula with predefined difficulty orderings in problems (Yang et al., 2024b; Team et al., 2025; Li et al., 2025), our approach dynamically assesses difficulty on the fly in each training step and applies difficulty filtering logic following the theoretical insights studied in §4. We describe the detailed process in Algorithm 1 and the high-level illustration of the algorithm in Figure 4 in Appendix A.
@@
First, we fill the batch $\mathcal{B}^{(t)}$ of the training step $t$ with filtered examples by measuring the success rate $p(\mathbf{x})$ (10) of each prompt $\mathbf{x}$ using sampled rollouts with size of $G$:
where $T_{\text{Low}}$ and $T_{\text{High}}$ are the predefined difficulty threshold. Here, the sample mean of $r_{acc}(\mathbf{x},\mathbf{y}_{i})$ is an unbiased estimate of $\mathbb{E}_{\mathbf{y}\sim\pi_{\theta_{t}}(\cdot|\mathbf{x})}\left[r_{\mathrm{acc}}(\mathbf{x},\mathbf{y})\right]$ with the sample size of $G$.
@@
While we showed that online difficulty filtering could maximize learnability in GRPO, naive filtering could result in inconsistent training batch size, leading to training instability and degraded performance (Li et al., 2022). For this reason, we ensure the fixed batch size to $|\mathcal{B}|=N$.
Rollouts for each prompt are sampled asynchronously and in parallel, enabling continuous batching of prompts and rollouts (Daniel et al., 2023; Kwon et al., 2023; Noukhovitch et al., 2025). Each prompt’s visit count, $\mathrm{vc}(\mathbf{x})$, is incremented after generating $G$ rollouts, ensuring it isn’t re-processed in the same iteration. Moreover, the active rollout process $\mathcal{P}_{\text{active}}$ is halted once the batch capacity is reached, allowing prompt training with the collected data. This sampling-based framework is compatible with Monte Carlo methods such as RLOO (Ahmadian et al., 2024) and VinePPO (Kazemnejad et al., 2024).
@@ 4.5 Difficulty filtering strategies
We mainly experiment two different difficulty filtering strategies, namely balanced difficulty filtering and skewed difficulty filtering:
@@
    Balanced difficulty filtering ($\mathrm{mean}(T_{\mathrm{High}},T_{\mathrm{Low}})=0.5$): We set the thresholds to be symmetric to the success rate of $0.5$: e.g., $T_{\mathrm{High}}=0.8$ and $T_{\mathrm{Low}}=0.2$.
@@
    Skewed difficulty filtering ($\mathrm{mean}(T_{\mathrm{High}},T_{\mathrm{Low}})\neq 0.5$): We set the thresholds asymmetrically, only filtering either easy or hard prompts: e.g., $T_{\mathrm{High}}=0.6$ and $T_{\mathrm{Low}}=0$.
By comparing the two strategies, we test if incorporating either side of extreme success rate cases can boost the performance of online difficulty filtering in GRPO, even though the theoretical learnability for either side has the same lower bound as analyzed in §4.3.
@@ 5.1 Experimental Setup
Before RORL experiments, we fine-tune Qwen2.5-3B base (Yang et al., 2024a) as a cold start, following the approach of Guo et al. (2025). Specifically, we curate 1.1K verified problem-solution pairs, with math problems sampled from NuminaMath (Li et al., 2024) and solutions distilled from DeepSeek-R1 (Guo et al., 2025).
Reinforcement learning.
For RORL, we employ GRPO on top of the SFT checkpoint. In each training step, the model generates 16 rollouts for 16 prompts (drawn from NuminaMath problems) and receives reward based on their correctness. We leave out 1,024 problems as a validation set. We also add a format reward and a language reward as in Guo et al. (2025). Additional training details for SFT and RORL are reported in the Appendix C.
5.2 Experimental design
Different strategies in online difficulty filtering.
Along with the plain GRPO without any prompt filtering, we test the online difficulty filtering with two different strategies introduced in §4.5: i.e., balanced and skewed filtering. For the balanced setting, we test $(T_{\mathrm{Low}},T_{\mathrm{High}})\in\left\{(0,1),(0.1,0.9),(0.2,0.8),(0.3,0.7),(0.4,0.6)\right\}$. For a skewed setting, we sweep $T_{\mathrm{Low}}\in\left\{0,0.2,0.4\right\}$ when $T_{\mathrm{High}}=1$ and $T_{\mathrm{High}}\in\left\{0.6,0.8,1\right\}$ when $T_{\mathrm{Low}}=0$.
Comparison against existing offline filtering methods.
We mainly compare two offline difficulty filtering methods with our approach: offline data curation (Yang et al., 2024b; Cui et al., 2025; Muennighoff et al., 2025; Ye et al., 2025) and offline scheduling (Team et al., 2025; Li et al., 2025). Offline data curation refers to the strategy that filters the problems by their difficulty before training, and offline scheduling additionally orders the training batches accordingly. For both offline strategies, we apply two settings, using Qwen2.5-7B-Instruct (Yang et al., 2024a) or our SFT model, as the difficulty proxies.
Evaluation Benchmarks.
We evaluate pass@1 on challenging math reasoning benchmarks: MATH500 (Hendrycks et al., 2021), AIME (Li et al., 2024), AMC (Li et al., 2024), MinervaMath (Lewkowycz et al., 2022), and OlympiadBench (He et al., 2024) (See Appendix D).
6 Results and Analysis
We first compare different online filtering strategies, balanced and skewed online filtering, in §6.1. Then, we compare with existing offline difficulty filtering methods, analyzing the impact of different difficulty assessment proxies in §6.2.
6.1 Online difficulty filtering strategies: balanced vs skewed filtering
@@
$1.00\leq p(\mathbf{x})<$ 1.0 55.4 3.3 22.8 19.8 19.8 24.2
$0.80\leq p(\mathbf{x})<$ 0.8 56.2 0.0 28.9 17.2 21.7 24.8
$0.60\leq p(\mathbf{x})<$ 0.6 56.2 3.3 26.5 21.3 21.6 25.8
@@
$10<p(\mathbf{x})<$ 1 60.8 3.3 31.3 18.0 27.3 27.3
$0.90.1<p(\mathbf{x})<$ 0.9 58.8 13.3 25.3 22.4 22.2 28.4
$0.80.2<p(\mathbf{x})<$ 0.8 62.2 10.0 30.1 20.5 26.3 29.8
$0.70.3<p(\mathbf{x})<$ 0.7 64.6 6.6 28.9 25.4 24.7 30.1
$0.60.4<p(\mathbf{x})<$ 0.6 60.2 6.6 32.8 25.0 24.9 29.9
Table 1: Five math reasoning benchmark evaluation results. “Minerva.” and “Olympiad.” refer to MinervaMath and OlympiadBench. “External” and “Initial” in offline filtering indicate using Qwen2.5-7B-Instruct and our SFT model as difficulty proxy for filtering. $p(\mathbf{x})$ (10) is the pass rate, the average correctness of rollouts. The highest and the second highest scores in each benchmark are highlighted with bold and underline, respectively.
@@
By tightening the pass rate threshold $(T_{\mathrm{Low}},T_{\mathrm{High}})$ for balanced filtering in Table 1, the average score of five benchmarks starts from 27.3% in $(0,1)$, gradually increasing until over 30% in $(0.3,0.7)$. Furthermore, simply removing examples in $(0,1)$ that do not contribute to learning in GRPO results in a slight improvement over the baseline, supporting the analysis in §4.2, i.e., $\hat{A}$ is zero for $(0,1)$. This result suggests that excluding ineffective examples improves both performance and training efficiency by focusing updates on meaningful data.
@@
While skewed filtering (“Skewed”) in Table 1 improves average performance by up to 5.7% over the SFT checkpoint, plain GRPO with 26.3% outperforms skewed filtering consistently in every threshold choice, which achieves around 24.9% to 25.9%. Overall, maximizing the expected learnability during GRPO training enhances learning in complex reasoning tasks. As discussed in §4.4, balanced filtering emerges as the best choice since it balances between penalizing and reinforcing diverse explorations.
6.2 Difficulty assessment proxy: offline vs online filtering
We apply Off-Diff with implementations from previous works (Yang et al., 2024a), with balanced threshold $(T_{\mathrm{Low}},T_{\mathrm{High}})=(0.2,0.8)$ following the results in §6.1.
Online difficulty filtering yields better learnability than offline methods.
While both offline curation (“Curation”) and offline scheduling (“Schedule”) in Table 1 show marginal improvements over plain GRPO with a maximum $2.1\%$ improvement, balanced online difficulty filtering consistently outperforms offline methods. Within offline methods, using an external difficulty assessment proxy (“External model” in Table 1) exceeded the case using the SFT checkpoint (“initial model”) on average, but with varying results by benchmark.
6.3 Analysis
Online filtering improves training efficiency.
Figure 2 illustrates the progression of the reward in the validation set, plotted against both the training steps (2(a)) and the training time on the wall clock (2(b)). As shown in Figure 2(a), models trained with balanced online difficulty filtering consistently outperform the plain GRPO ($0\leq p(\mathbf{x})\leq 1$) in fewer training steps. This suggests that by filtering out less informative examples, the average learnability within each batch increases, allowing faster learning progress. Interestingly, Figure 2(b) shows that this benefit carries over even when measured by wall-clock time by exceeding plain GRPO’s maximum reward in less training time. However, we also observe that overly aggressive filtering, such as in the case of the $0.60.4<p(\mathbf{x})<$ 0.6 setting, can require significantly more rollouts to fill a batch, leading to longer training times overall. These results suggest that online filtering can enable more efficient learning even in real-world settings, as long as overly aggressive filtering is avoided.

### v2

@@ 3.2 Prompt-level learnability: theoretical analysis
$\begin{split}&\mathbb{E}_{y\sim\pi_{\mathrm{init}}(\cdot|x)}~\left[\log\frac{\pi^{*}(y|x)}{\pi_{\mathrm{init}}(y|x)}\right]\\ &=\frac{p(x)}{\beta}-\log~\Big((1-p(x))+p(x)e^{1/\beta}\Big),\end{split}$ (13)
and its negative corresponds to the reverse KL divergence $\mathbb{D}_{\mathrm{KL}}\left(\pi_{\mathrm{init}}(\cdot|x)\,\|\,\pi^{*}(\cdot|x)\right)$ defined in equation 9.
Proposition 3.1 (Variance-controlled separation and degeneracy).
For any prompt $x$ and temperature $\beta>0$, the reverse KL divergence between the initial and optimal policies, $\pi_{\mathrm{init}}$ and $\pi^{*}$, satisfies
$\mathbb{D}_{\mathrm{KL}}~\big(\pi_{\mathrm{init}}(\cdot|x)\|\pi^{*}(\cdot|x)\big)\;\geq\;\frac{p(x)\big(1-p(x)\big)}{2\beta^{2}},$ (14)
with the bound maximized at $p(x)=\tfrac{1}{2}$. In particular, for absolute-hard or absolute-easy prompts where $p(x)\in\{0,1\}$, the divergence vanishes and $\pi_{\mathrm{init}}$ is already optimal:
$\mathbb{D}_{\mathrm{KL}}~\big(\pi_{\mathrm{init}}(\cdot|x)\,\|\,\pi^{*}(\cdot|x)\big)=0.$ (15)
Proof sketch. Starting from equation 13 and expanding $\log~\big((1-p)+pe^{1/\beta}\big)$ with $\log(1+\epsilon)\geq\epsilon-\tfrac{\epsilon^{2}}{2}$ for $\epsilon=p(x)(1/\beta+1/(2\beta^{2}))$, we obtain
$\mathbb{E}_{y\sim\pi_{\mathrm{init}}}\left[\log\frac{\pi^{*}(y|x)}{\pi_{\mathrm{init}}(y|x)}\right]\leq-\frac{p(x)\big(1-p(x)\big)}{2\beta^{2}}.$ (16)
Since the reverse KL divergence is the negative of this expectation equation 9, the inequality in equation 14 follows. When $p(x)\in\{0,1\}$, both the expected reward and the soft value function coincide ($V^{*}(x)=r_{\mathrm{acc}}(x,y)$), yielding zero divergence. The complete proof is provided in Appendix C. $\square$
Remark 3.2 (Learnability in absolute prompts: No learning signal at extremes).
For absolute-hard or absolute-easy prompts ($p(x)\in\{0,1\}$), the advantage term in GRPO equation 4 becomes zero for all rollouts, indicating that such prompts provide no gradient signal during policy optimization.
Remark 3.3 (Learnability in soft prompts: Maximal learnability band).
The lower bound in equation 14 is proportional to the Bernoulli variance p(x)(1p(x))p(x)(1-p(x)), which attains its maximum at $p(x)=0.5$. Hence, prompts whose pass rates lie in the intermediate region $\epsilon\leq p(x)\leq 1-\epsilon$ provide the strongest learning signal, while soft-hard ($p(x)\approx\epsilon$) and soft-easy ($p(x)\approx 1-\epsilon$) prompts contribute only marginally.
Proposition 3.1 unifies both extreme and intermediate cases of prompt learnability. When $p(x)$ approaches the extremes, the reverse KL divergence $\mathbb{D}_{\mathrm{KL}}~\big(\pi_{\mathrm{init}}\|\pi^{*}\big)$ tends to zero, implying no separation between policies and thus no effective update. Conversely, when $p(x)\approx 0.5$, the reward variance, i.e., learnability, is maximized. These insights directly motivate the filtering criterion in our online curriculum, which selectively retains prompts within the intermediate difficulty band to ensure the highest learning efficiency.
@@
1:Initial policy model $\pi_{\mathrm{init}}$; Reward $r$; Prompts queue $\mathcal{Q}$; Pass rate thresholds $T_{\text{Low}},T_{\text{High}}$; Batch size NN; Group size $G$; $r_{\text{acc}}$ equation 5; Visit count $\mathrm{vc}(\mathbf{x})$.
[…]
@@
21:   Compute $\hat{A}_{i}$ for $\mathbf{y}_{i}$ in $\mathcal{B}^{(t)}$ through group relative advantage estimation equation 4.
@@ 3.3 Method: online difficulty filtering with fixed batch size
We analyze an online difficulty filtering approach that ensures a fixed batch size throughout training for a reasoning-oriented agent. Unlike static curricula with predefined difficulty orderings in problems (Yang et al., 2024c; Team et al., 2025; Li et al., 2025), our approach dynamically assesses difficulty on the fly in each training step and applies difficulty filtering logic following the theoretical insights studied in §3. We describe the detailed process in Algorithm 1 and the high-level illustration of the algorithm in Figure 4 in Appendix B.
@@
First, we fill the batch $\mathcal{B}^{(t)}$ of the training step $t$ with filtered examples by measuring the success rate $p(\mathbf{x})$, equation 11, of each prompt $\mathbf{x}$ using sampled rollouts with size of $G$ as in equation 25. With the predefined difficulty threshold $T_{\text{Low}}$ and $T_{\text{High}}$, we asynchronously filter and fill the batch to meet the fixed batch size.
@@
While we showed that online difficulty filtering could maximize learnability in GRPO, naive filtering could result in inconsistent training batch size, leading to training instability and degraded performance (Li et al., 2022). For this reason, we ensure the fixed batch size to $|\mathcal{B}|=N$ for the batch $\mathcal{B}^{(t)}$ at the training step $t$ as in equation 25. We extend technical details in Appendix B.
@@ 3.4 Difficulty filtering strategies
Based on the literature study in Appendix A, we experiment two different difficulty filtering strategies, balanced and skewed difficulty filtering:
@@
    Balanced difficulty filtering: We set the thresholds to be symmetric to the success rate of $0.5$: e.g., $T_{\mathrm{High}}=0.8$ and $T_{\mathrm{Low}}=0.2$.
@@
    Skewed difficulty filtering: We set asymmetric thresholds, only filtering either easy or hard prompts: e.g., $T_{\mathrm{High}}=0.6$ and $T_{\mathrm{Low}}=0$.
We test if incorporating either side of extreme pass rate cases can boost the performance of online difficulty filtering, even though the learnability for either side has the same lower bound as in §3.2.
@@ 4.1 Experimental Setup
Before reinforcement learning with verifiable rewards (RLVR) experiments, we fine-tune Qwen2.5-3B base (Yang et al., 2024b) as a cold start, following Guo et al. (2025). Specifically, we curate 1.1K verified problem-solution pairs, with math problems sampled from NuminaMath (Li et al., 2024) and solutions distilled from DeepSeek-R1 (Guo et al., 2025).
Reinforcement learning with verifiable rewards.
For RLVR, we employ GRPO on top of the SFT checkpoint. In each training step, the model generates 16 rollouts for 16 prompts drawn from NuminaMath problems and receives a reward based on their correctness. We leave out 1,024 problems as a validation set. We also add a format reward and a language reward as in Guo et al. (2025). Additional training details are reported in the Appendix D.
Table 1: Five math reasoning benchmark evaluation results with Qwen2.5-3B. “Minerva.” and “Olympiad.” refer to MinervaMath and OlympiadBench. “External” and “Initial” in offline filtering indicate using Qwen2.5-7B-Instruct and our SFT model as a difficulty proxy for filtering. $p(\mathbf{x})$ equation 11 is the pass rate, the average correctness of rollouts. The highest and the second highest scores in each benchmark are highlighted with bold and underline.
@@
$0\leq p(\mathbf{x})<1.0$ 55.4 3.3 22.8 19.8 19.8 24.2
$0\leq p(\mathbf{x})<0.8$ 56.2 0.0 28.9 17.2 21.7 24.8
$0\leq p(\mathbf{x})<0.6$ 56.2 3.3 26.5 21.3 21.6 25.8
@@
$0<p(\mathbf{x})<1$ 60.8 3.3 31.3 18.0 27.3 27.3
$0.1<p(\mathbf{x})<0.9$ 58.8 13.3 25.3 22.4 22.2 28.4
$0.2<p(\mathbf{x})<0.8$ 62.2 10.0 30.1 20.5 26.3 29.8
$0.3<p(\mathbf{x})<0.7$ 64.6 6.6 28.9 25.4 24.7 30.1
$0.4<p(\mathbf{x})<0.6$ 60.2 6.6 32.8 25.0 24.9 29.9
4.2 Experimental design
Different strategies in online difficulty filtering.
Along with the plain GRPO without any prompt filtering, we test the online difficulty filtering with two different strategies introduced in §3.4: i.e., balanced and skewed filtering. For the balanced setting, we test $(T_{\mathrm{Low}},T_{\mathrm{High}})\in\left\{(0,1),(0.1,0.9),(0.2,0.8),(0.3,0.7),(0.4,0.6)\right\}$. For a skewed setting, we sweep $T_{\mathrm{Low}}$ in $\left\{0,0.2,0.4\right\}$ when $T_{\mathrm{High}}=1$ and $T_{\mathrm{High}}$ in $\left\{0.6,0.8,1\right\}$ when $T_{\mathrm{Low}}=0$.
Comparison against existing offline filtering methods.
We mainly compare two offline difficulty filtering methods with our approach: offline data curation (Yang et al., 2024c; Cui et al., 2025; Muennighoff et al., 2025; Ye et al., 2025) and offline scheduling (Team et al., 2025; Li et al., 2025). Offline data curation refers to the strategy that filters the problems by their difficulty before training, and offline scheduling additionally orders the training batches accordingly. For both offline strategies, we used Qwen2.5-7B-Instruct (Yang et al., 2024b) or our SFT model as the difficulty proxies.
Evaluation Benchmarks.
We evaluate pass@1 across math reasoning benchmarks of varying difficulty levels: MATH500 (Hendrycks et al., 2021), AIME (Li et al., 2024), AMC (Li et al., 2024), MinervaMath (Lewkowycz et al., 2022), and OlympiadBench (He et al., 2024) (See Appendix E).
5 Results
We first compare different online filtering strategies in §5.1 and expand to existing offline difficulty filtering methods in §5.2.
5.1 Balanced and skewed filtering
@@
Table 2: Five math reasoning benchmark evaluations with Qwen2.5-7B. The notations follow that of Table 1.
Method Difficulty Filter MATH500 AIME AMC Minerva Olympiad Avg.
SFT 72.6 12.1 34.9 32.0 35.1 37.3
GRPO w/ Online Filtering (Ours) Plain
$0\leq p(\mathbf{x})\leq 1$ 75.0 12.3 42.2 32.7 35.7 39.6
Balanced
$0<p(\mathbf{x})<1$ 75.0 13.1 41.0 33.5 36.9 39.9
$0.3<p(\mathbf{x})<0.7$ 75.8 15.0 47.0 33.8 37.6 41.8
@@
By tightening the pass rate threshold $(T_{\mathrm{Low}},T_{\mathrm{High}})$ for balanced filtering in Table 1, the average score of five benchmarks starts from 27.3% in $(0,1)$, gradually increasing until over 30% in $(0.3,0.7)$. Furthermore, simply removing examples in $(0,1)$ that do not contribute to learning in GRPO results in a slight improvement over the baseline, aligning with Remark 3.2, i.e., $\hat{A}$ is zero for $(0,1)$. This result suggests that excluding ineffective examples improves both performance and training efficiency by focusing updates on meaningful data. These observations are further supported by the difficulty-level analysis using MATH500 that provides five different levels in Appendix F, which shows consistent gains across different levels.
@@
While skewed filtering (“Skewed”) in Table 1 improves average performance up to 5.7% over the SFT checkpoint, plain GRPO with 26.3% outperforms skewed filtering consistently in every threshold choice, which achieves around 24.9% to 25.9%. Overall, maximizing the expected learnability in GRPO enhances learning in complex reasoning tasks. As discussed in §3.3, balanced filtering emerges as the best choice since it balances between penalizing and reinforcing diverse explorations.
5.2 Offline and online filtering
We apply the offline difficulty filtering with implementations from previous works (Yang et al., 2024b), with balanced threshold $(T_{\mathrm{Low}},T_{\mathrm{High}})=(0.2,0.8)$ following the results in §5.1. While both offline curation (“Curation”) and offline scheduling (“Schedule”) in Table 1 show marginal improvements over plain GRPO with a maximum $2.1\%$ improvement, balanced online difficulty filtering consistently outperforms offline methods. Within offline methods, using an external difficulty assessment proxy (“External model”) exceeded the case using the SFT checkpoint (“initial model”) on average, but with varying results by benchmark.
5.3 Scalability of balanced online filtering
We adopt 7B scale model within the same Qwen2.5 family to confirm the scalability of the proposed method. In Table 2, stricter filtering thresholds ($0.3<p(\mathbf{x})<0.7$) yield the strongest performance with 3% and 5% increase in AIME and AMC, respectively. Overall, the ascending trend in Table 2 aligns with the 3B cases, demonstrating the scalability of online difficulty filtering.
6 Analysis
6.1 Learning dynamics analysis
Figure 2: Perceived difficulty per batch curated through balanced online filtering. Defining “difficulty” as $1-p(\mathbf{x})$, a greater difficulty implies lower sample accuracy.